number of ways of distributing n identical objects in to r groups when empty group not allowed?

Grade:

1 Answers

Abhishek Kumar
8 years ago
We have to calculate,
nC1 + nC2 + nC3 + ….…. + nCr
Now, (1+x)r = 2r ; when x =1
So,
nC0 + nC1 + nC2 + nC3 + ….…. + nCr = 2r
nC1 + nC2 + nC3 + ….…. + nCr = 2r -1